# Repositioning, Ride-matching, and Abandonment in On-demand Ride-hailing Platforms: A Mean Field Game Approach

### v1

@@ 3.3. Formulas of Waiting Time and Pick-up Time
In this section we establish formulas for the average waiting time and pick-up time as claimed in Assumption 1 based on the customer abandonment behavior defined in Section 3.1 and the platform’s dispatching algorithm defined in Section 3.2.
@@
Let each region $i$’s area be denoted by $A_{i}$. We introduce two matching radii in region $i$, $R_{i,c},R_{i,d}\in(0,\sqrt{A_{i}/\pi})$, where $\sqrt{A_{i}/\pi}$ is the approximate radius of region $i$. We denote the average vehicle speed by $v_{i}$. Let $\mu_{i}^{c}$ and $\mu_{i}^{d}$ denote the average densities of waiting customers and waiting drivers respectively across the region $i$. Recall that $x_{ii}$ denotes rate of drivers that become available and choose to serve region $i$ per unit area, and $b_{i}$, $\theta_{i}$ the arrival rate, abandonment rate of customers per unit area, respectively.
$(1-\frac{\pi R_{i,c}^{2}}{A_{i}})^{N_{i}^{c}}.$
Similarly, given the matching radius $R_{i,d}$, if there are $N_{i}^{d}$ waiting drivers in region $i$, a new arrival customer is not matched with any of these waiting customers with probability
$(1-\frac{\pi R_{i,d}^{2}}{A_{i}})^{N_{i}^{d}}.$
(15) $x_{ii}(1-\frac{\pi R_{i,c}^{2}}{A_{i}})^{\mu_{i}^{c}A_{i}}+b_{i}(1-\frac{\pi R_{i,d}^{2}}{A_{i}})^{\mu_{i}^{d}A_{i}}=b_{i}$
(16) $\mu_{i}^{c}(x_{ii})=\frac{1}{\theta}(b_{i}-x_{ii}).$
As a result, by (15), the unknown variable $\mu_{i}^{d}$ can also be expressed a as a function of $x_{ii}$,
Next, we derive the formula for expected pick-up time, which is amount to finding the expected pick-up distance. We compute this in two distinct cases:
    $1-(1-\frac{\pi R_{i,c}^{2}}{A_{i}})^{\mu_{i}^{c}A_{i}}.$
    $1-(1-\frac{\pi r^{2}}{A_{i}})^{\mu_{i}^{c}A_{i}}$
    $(1-\frac{\pi R_{i,c}^{2}}{A_{i}})^{\mu_{i}^{c}A_{i}}.$
    $1-(1-\frac{\pi R_{i,d}^{2}}{A_{i}})^{\mu_{i}^{d}A_{i}}.$
    $1-(1-\frac{\pi r^{2}}{A_{i}})^{\mu_{i}^{d}A_{i}}$
In summary, the expected pick-up time is a function of $x_{ii}$ as follows,
where $\mu_{i}^{c}$ and $\mu_{i}^{d}$ are functions of $x_{ii}$ as in (16) and (17) respectively.
$x_{ii}e^{-\mu_{i}^{c}\pi R_{i,c}^{2}}+b_{i}e^{-\mu_{i}^{d}\pi R_{i,d}^{2}}=b_{i}.$
(21) $\begin{cases}w_{i}^{d}(x_{ii})&=\frac{-\log(1-\frac{x_{ii}}{b_{i}}e^{-\frac{b_{i}-x_{ii}}{\theta_{i}}\pi R_{i,c}^{2}})}{x_{ii}\pi R_{i,d}^{2}}\\ \tau_{i}(x_{ii})&=\frac{1}{v_{i}}\int_{0}^{R_{i,c}}e^{-\mu_{i}^{c}(x_{ii})\pi r^{2}}-e^{-\mu_{i}^{c}(x_{ii})\pi R_{i,c}^{2}}\mathrm{d}r+\frac{b_{i}}{v_{i}x_{ii}}\int_{0}^{R_{i,d}}e^{-\mu_{i}^{d}(x_{ii})\pi r^{2}}-e^{-\mu_{i}^{d}(x_{ii})\pi R_{i,d}^{2}}\mathrm{d}r\end{cases}$
Note that the left hand side of (20) is approximately $R_{i,\cdot}^{2}$ when $R_{i,\cdot}<<\sqrt{A_{i}}$ as $\log(1-y)\approx-y$ when $y$ approaches 0 and goes to $\infty$ as $\pi R_{i,\cdot}^{2}\rightarrow A_{i}$. Hence, when calculating the pick-up time we do not multiply the change of variable factor when computing the integral. Numerical results show that this is still a good estimation of the original pick-up time even when the original radius $R_{i}$ is large. There is an alternative point of view which yields the same approximate formula in (21): assume each region $i$ is an unbounded plane with matching radii $R_{i,c},R_{i,d}\in(0,\infty)$. In this unbounded plane, the locations of waiting customers and drivers follow spatial Poisson processes, where $\mu_{i}^{c}$ and $\mu_{i}^{d}$ are the waiting customer and waiting driver intensity parameters, respectively. In fact, the change of variable in equation (20) defines a map from the original finite region to an infinite region. Although this transformation distorts the Euclidean metric in the original region, the distortion is negligible when the matching radius is sufficiently small. For large matching radii, while the distortion cannot be entirely ignored, the introduced approximation error remains acceptable according to numerical results.
3.4. Analysis of Equilibria in One Region: Wild Goose Chase
@@
In this section, we show that the original NN dispatch algorithm, where the matching radius $R_{i,\cdot}$ is set to $\sqrt{A_{i}/\pi}$ (or $\infty$ if the approximated formulas (21) are used), can still lead to multiple equilibria even when there is only one region. To simplify notation, we omit the subscript $i$ since there is only one region. As there is no driver repositioning, $x$ denotes the driver supply rate. When the two matching radii are $\sqrt{A/\pi}$, the equation (15) becomes,
$x\cdot 0^{\mu^{c}}+b\cdot 0^{\mu^{d}}=b$
@@
which seems to be meaningless. But if we define the indeterminate limit [^2] Note that $\lim\limits_{x\rightarrow 0}\lim\limits_{y\rightarrow 0}x^{y}=1$ and $\lim\limits_{y\rightarrow 0}\lim\limits_{x\rightarrow 0}x^{y}=0$. $0^{0}$ to be $1$, we obtain two possible solutions,
@@
These solutions are validated through an intuitive argument: it is impossible for both the number of waiting customers and the number of waiting drivers to be positive when the matching radius is infinite, as they would be matched with each other. Consequently, there are two distinct types of equilibriua:
@@
Before presenting the equilibrium results, we must first discuss how to calculate the pick-up time using Equation (19) or the approximate formula in Equation (21) under this extreme case.
Approximate Pick-up Time. When $R_{c}=R_{d}=\sqrt{A/\pi}$ (or $\infty$), our formulae for $\tau_{i}$ in (19) and (21) can be still used to estimate the pick-up time. In fact, both formulas give the same asymptotical result. We assume $\mu^{c}=0,\mu^{d}>0$. In this case, $x=b$ as $\mu^{c}=0$. Applying (19), the average pick-up time is estimated by
We analyze the two types of equilibria as follows.
Case 1: Oversupply Equilibrium. In this case, there is an oversupply of drivers, and hence a positive density of waiting drivers, i.e., $\mu^{c}=0,\mu^{d}>0$. Given the pick-up time (22), the conversation law of mass (3) gives,
$\mu^{d}+\frac{b}{2v\sqrt{\mu^{d}}}+bt=m$
which has one solution of equilibrium when $m=\tfrac{1}{3}\left(\tfrac{b^{2}}{16v^{2}}\right)^{\frac{1}{3}}+bt$, and two solutions of equilibria when $m>\tfrac{1}{3}\left(\tfrac{b^{2}}{16v^{2}}\right)^{\frac{1}{3}}+bt.$ This type of equilibria have the same driver supply rate. We denote the equilibrium driver supply rate in this case by $x^{o}$ and then $x^{o}=b$ since every customer is served in this oversupply regime.
Case 2: Undersupply Equilibrium. In this case, there is an oversupply of drivers, and hence a positive density of waiting customers, i.e., $\mu^{c}>0,\mu^{d}=0$. Then $w^{d}=0$ as $\mu^{d}=0$. The pick-up time is given by a square root law similar to (22), and the conversation law of mass (3) becomes,
(23) $\frac{x}{2v\sqrt{\mu^{c}}}+xt=\frac{x}{2v}\sqrt{\frac{\theta}{b-x}}+xt=m$
which has at least one solution of equilibrium as long as $m>0$. We denote the equilibrium service rate in this case by $x^{u}$ where $x^{u}$ solves equation (23). Here, $x^{u}<b$, which is a “bad” equilibrium compared to the oversupply regime. The equation (23) may admit solutions with extremely small driver supply rates and large pick-up times.
@@
Infinity Price of Anarchy. The undersupply equilibrium can be arbitrarily bad compared to the oversupply equilibrium, i.e., $x^{u}<<b$. Specifically, when the mass is $m=\tfrac{1}{3}\left(\tfrac{b^{2}}{16v^{2}}\right)^{\frac{1}{3}}+bt$, there are exactly two equilibria The two equilibria are $x^{o}=b$ and $x^{u}$ where $\frac{x^{u}}{b}$ satisfies
@@
which approaches 00 if $\theta b^{3}=1$ and $b\rightarrow 0$. Therefore, the “bad” equilibrium ($x^{u}$) can be arbitrarily worse than the “good” equilibrium ($x^{o}$). Furthermore, an infinity price of anarchy occurs under the following conditions: 1) the region size is sufficiently large ($A\rightarrow\infty$); 2) the customer arrival rate is extremely low ($b\rightarrow 0$); and 3) the customers exhibit extreme impatience ($\theta\rightarrow\infty$). This extreme scenario is primarily theoretical and rarely observed in practice. However, it highlights a critical insight: if the dispatch algorithm is not properly designed, the system’s efficiency loss can be catastrophic.
Wild Goose Chase. The WGC equilibrium exists in the undersupply regime (characterized by the “bad” equilibrium $x^{u}$), where drivers are assigned to pick up customers located far away. In this regime, the system experiences a driver shortage because: 1) high average pick-up times force drivers to travel long distances to reach customers; 2) wasted time en route further reduces driver availability, exacerbating the undersupply. This creates a vicious cycle: longer travel times lead to fewer drivers being available to respond to new requests, worsening the imbalance between supply and demand.
@@
Figure 2. The driver density (RHS of (24)) as a function of the driver supply rate $x$. We set the matching radius $R_{c}$ and $R_{d}$ to be same. In this figure, from left to right, the value of matching radius is 0.5 km, 1 km, 2 km, and 3 km.
Impact of Matching Radius. The use of fixed matching radius with $R_{c},R_{d}<\sqrt{A/\pi}$, does not completely eliminate multiple equilibria or prevent the WGC equilibrium. In the one region case. we can compute the equilibria by employing the conversation law of mass in (3), i.e.,
@@
where $w^{d}(x)$ and $\tau(x)$ are given in (18) and (19) . Since $\tau(x)$ is non-monotonic, multiple equilibria can arise for specific parameter choices.
@@
Extension to Multiple Region. When multiple regions are considered, the average pick-up time in each region is a non-monotonic function of the local driver supply rate, leading to the potential for even more equilibria. In the single-region case, there can be three equilibria. As the number of regions increases, the number of equilibria grows combinatorially. We omit the detailed analysis here due to its complexity, but this can be verified using a toy model, such as a spoke-and-hub network.

### v2

@@ 3.3. Formulas for Waiting Times and Pick-up Times
In this section we establish formulas for the average waiting time and pick-up time as functions of the driver supply rate, as required (Assumption 1) for applying Theorem 2.3 to identify equilibria through an optimization problem, based on the customer abandonment behavior defined in Section 3.1 and the platform’s dispatching algorithm defined in Section 3.2.
@@
Let each region $i$’s area be denoted by $a_{i}$. We introduce two matching radii in region $i$, $R_{i,c},R_{i,d}\in(0,\sqrt{a_{i}/\pi})$, where $\sqrt{a_{i}/\pi}$ is the approximate maximum radius of region $i$. We denote the average vehicle speed by $v_{i}$. Let $\mu_{i}^{c}$ and $\mu_{i}^{d}$ denote the average densities of waiting customers and waiting drivers respectively across the region $i$. Recall that $x_{ii}$ denotes the rate of drivers that become available and choose to serve region $i$ per unit area, and $b_{i}$, $\theta_{i}$ the arrival rate and abandonment rate of customers per unit area, respectively.
We assume the locations of waiting customers and waiting drivers are drawn at any moment from independent uniform distributions across region $i$. This assumption is justified by the randomness of customers’ origins and destinations, as well as the mobility of the vehicles. We can now derive the formula for the driver’s average waiting time. A critical observation is that the rate at which drivers enter the waiting state is equal to the rate at which they leave it. The entry rate is exactly the rate at which drivers become idle after failing to be matched with customers, while the exit rate is the rate at which the waiting drivers are matched with arriving customers. Using this key insight, we first establish a relationship between $\mu_{i}^{c}$ and $\mu_{i}^{d}$:
(15) $x_{ii}\left(1-\frac{\pi R_{i,c}^{2}}{a_{i}}\right)^{\mu_{i}^{c}a_{i}}=b_{i}\left(1-\left(1-\frac{\pi R_{i,d}^{2}}{a_{i}}\right)^{\mu_{i}^{d}a_{i}}\right)\,,$
where the left-hand side represents the approximate average rate at which drivers enter the waiting state, while the right-hand side represents the approximate average rate at which the waiting drivers are matched with arriving customers. With the relations (14) and (15), as well as Little’s law, $\mu_{i}^{d}=x_{ii}w_{i}^{d}$, we are able to derive the formulas for driver’s average waiting time and pick-up time, whose detailed derivation are provided in Appendix A.
Proposition 3.1.
Assume the platform adopts the two-matching-radius nearest-neighbor dispatch with radii $R_{i,c}$ and $R_{i,d}$ in each region $i\in S$. Then, the driver’s mean waiting time and mean pickup time can be expressed as explicit functions of the driver supply rate $x_{ii}$ as follows,
(16) $w_{i}^{d}(x_{ii})$ $=\frac{\log(1-\frac{x_{ii}}{b_{i}}(1-\frac{\pi R_{i,c}^{2}}{a_{i}})^{\frac{b_{i}-x_{ii}}{\theta_{i}}a_{i}})}{x_{ii}a_{i}\log(1-\frac{\pi R_{i,d}^{2}}{a_{i}})}\,,$
(17) $\tau_{i}(x_{ii})$ $=\frac{1}{v_{i}}\int_{0}^{R_{i,c}}(1-\frac{\pi r^{2}}{a_{i}})^{\mu_{i}^{c}a_{i}}-(1-\frac{\pi R_{i,c}^{2}}{a_{i}})^{\mu_{i}^{c}a_{i}}\mathrm{d}r+\frac{b_{i}}{v_{i}x_{ii}}\int_{0}^{R_{i,d}}(1-\frac{\pi r^{2}}{a_{i}})^{\mu_{i}^{d}a_{i}}-(1-\frac{\pi R_{i,d}^{2}}{a_{i}})^{\mu_{i}^{d}a_{i}}\mathrm{d}r.$
Proposition 3.1 validates Assumption 1. More specifically, Proposition 3.1 states that the driver’s mean waiting time and mean pickup time can indeed be expressed as explicit functions of the driver supply rate as required by Assumption 1.(1) . Furthermore, one can check that as $x_{ii}\rightarrow b_{i}$, $w_{i}^{d}(x_{ii})$ in (16) goes to $\infty$, which satisfies Assumption 1.(2) . As a result, by Theorem 2.3, the equilibria of game $\mathcal{G}$ under our two-matching-radius nearest-neighbor dispatch algorithm are exactly the KKT points of the optimization problem (10) with $\tau_{i}$ and $w_{i}^{d}$ given by (16) and (17) respectively.
To approximate the formulas (16) and (17), we apply the following change of variable,
(18) $-\frac{a_{i}}{\pi}\log(1-\frac{\pi R_{i,\cdot}^{2}}{a_{i}})\mapsto R_{i,\cdot}^{2}.$
where with a slight abuse of notation we keep the same variable for the new radius. In fact, the left-hand side of (18) is approximately $R_{i,\cdot}^{2}$ when $R_{i,\cdot}<<\sqrt{a_{i}}$ as $\log(1-y)\approx-y$ when $y$ approaches 0. Then, the drivers’ waiting times and pick-up times can be approximated as:
(19) $\begin{cases}w_{i}^{d}(x_{ii})&=\frac{-\log(1-\frac{x_{ii}}{b_{i}}e^{-\frac{b_{i}-x_{ii}}{\theta_{i}}\pi R_{i,c}^{2}})}{x_{ii}\pi R_{i,d}^{2}}\,,\\ \tau_{i}(x_{ii})&=\frac{1}{v_{i}}\int_{0}^{R_{i,c}}e^{-\mu_{i}^{c}(x_{ii})\pi r^{2}}-e^{-\mu_{i}^{c}(x_{ii})\pi R_{i,c}^{2}}\mathrm{d}r+\frac{b_{i}}{v_{i}x_{ii}}\int_{0}^{R_{i,d}}e^{-\mu_{i}^{d}(x_{ii})\pi r^{2}}-e^{-\mu_{i}^{d}(x_{ii})\pi R_{i,d}^{2}}\mathrm{d}r\,.\end{cases}$
With these formulas, we are able to analyze the equilibria of the game in the next section.
3.4. Equilibria Analysis
@@
In this section, we show that the original NN dispatch algorithm, where the matching radius $R_{i,\cdot}$ is set to the maximal value $\sqrt{a_{i}/\pi}$, i.e., the approximated radius of region $i$, (or $\infty$ if the approximated formulas (19) are used), can lead to multiple equilibria even when there is only one region. Furthermore, we examine the influence of the matching radius in the one-region case and discuss how these findings could be extended to multi-region scenarios.
We first analyze the equilibrium when there is only one region. To simplify notation, we omit the subscript $i$. As there is no driver repositioning, we use $x$ to denote the driver supply rate. When the two matching radii are $\sqrt{a/\pi}$, the equation (15) becomes,
$x\cdot 0^{\mu^{c}}+b\cdot 0^{\mu^{d}}=b,$
@@
which seems to be meaningless. But if we define the indeterminate limit [^3] Note that $\lim\limits_{x\rightarrow 0}\lim\limits_{y\rightarrow 0}x^{y}=1$ and $\lim\limits_{y\rightarrow 0}\lim\limits_{x\rightarrow 0}x^{y}=0$. $0^{0}$ to be $1$, we obtain two possible solutions for $\mu^{c},\mu^{d}$,
@@
These solutions are validated through an intuitive argument: it is impossible for both the number of waiting customers and the number of waiting drivers to be positive when the matching radius is infinite, as they would be matched with each other. Consequently, there are two distinct types of equilibria:
@@
We next analyze the two types of equilibria. As either $\mu^{c}=0,\mu^{d}>0\text{ or }\mu^{c}>0,\mu^{d}=0$, we assume the following form of square root law to estimate the average pick-up time,
(20) $\tau=\begin{cases}\frac{1}{2v\sqrt{\mu^{c}}}&\text{ if }\mu^{d}=0,\\ \frac{1}{2v\sqrt{\mu^{d}}}&\text{ if }\mu^{c}=0.\end{cases}$
Similar forms of square root law are also used in related literature (Besbes et al., 2022; Wang et al., 2022). In Appendix B, we demonstrate that the formulas in (17) and (19) both lead to (20) when $R_{i,\cdot}\rightarrow\sqrt{a_{i}/\pi}$ and $R_{i,\cdot}\rightarrow\infty$ respectively through an asymptotical analysis. In the following discussion, we use $x^{o}$ and $x^{u}$ to denote the equilibrium driver supply rates for oversupply type and undersupply type respectively.
Type 1: Oversupply Equilibrium. In this type of equilibrium, there is an oversupply of drivers, resulting in a positive density of waiting drivers, i.e., $\mu^{c}=0$ and $\mu^{d}>0$. The driver supply rate in this equilibrium is always equal to $b$, i.e., $x^{o}=b$, as every customer is served in the oversupply equilibrium. The only unknown variable for this type of equilibrium is the density of waiting drivers, $\mu^{d}$. To find the unknown $\mu^{d}$, we apply the pick-up time formula in (20) and solve the conservation law of mass equation as follows,
(21) $\mu^{d}+\frac{b}{2v\sqrt{\mu^{d}}}+bt=m\,,$
which has one solution when $m=\tfrac{1}{3}\left(\tfrac{b^{2}}{16v^{2}}\right)^{\frac{1}{3}}+bt,$ and two solutions when $m>\tfrac{1}{3}\left(\tfrac{b^{2}}{16v^{2}}\right)^{\frac{1}{3}}+bt.$
Type 2: Undersupply Equilibrium. In this type of equilibrium, there is an undersupply of drivers, resulting in a positive density of waiting customers, i.e., $\mu^{c}>0$ and $\mu^{d}=0$. According to Little’s law, $w^{d}=\mu^{d}/x=0$, as there are no waiting drivers. In this case, the unknown variable is the driver supply rate, $x$. To find the unknown $x$, we use the pick-up time formula in (20) and solve the conservation law of mass equation $\tfrac{x}{2v\sqrt{\mu^{c}}}+xt=m$. By substituting (14), the equation becomes
(22) $\frac{x}{2v}\sqrt{\frac{\theta}{b-x}}+xt=m,$
which has exactly one solution as long as $m>0$. The equilibrium driver supply rate $x^{u}$ satisfies equation (22) and $x^{u}<b$, representing an inefficient equilibrium compared to the oversupply one.
We summarize all the possible equilibria results as follows.
Proposition 3.2.
Assume there is a single region and the platform adopts the NN dispatch algorithm with no restriction on the matching radius. There can be multiple equilibria depending on the size of driver density, $m$:
  • 1)
    When $0<m<\tfrac{1}{3}\left(\tfrac{b^{2}}{16v^{2}}\right)^{\frac{1}{3}}+bt$, there is only one equilibrium which is an undersupply one with driver supply rate $x^{c}$ the unique solution to (22) and zero waiting driver density.
  • 2)
    When $m=\tfrac{1}{3}\left(\tfrac{b^{2}}{16v^{2}}\right)^{\frac{1}{3}}+bt$, there are two equilibria: one oversupply equilibrium with driver supply rate $x^{o}=b$ and waiting driver density $\mu^{d}$ the unique solution to (21), and one undersupply equilibrium with driver supply rate $x^{u}$ the unique solution to (22), and zero waiting driver density .
  • 3)
    When $m>\tfrac{1}{3}\left(\tfrac{b^{2}}{16v^{2}}\right)^{\frac{1}{3}}+bt$, there are three equilibria: two oversupply equilibrium with the same driver supply rate $x^{o}=b$ and two different waiting driver densities given by solutions to (21) respectively, and one undersupply equilibrium with driver supply rate $x^{u}$ the unique solution to (22) and zero waiting driver density.
Proposition 3.2 demonstrates that there can be multiple equilibria under the NN dispatch algorithm with no restriction on the matching radius even in a single region. In case 2) and 3) of Proposition 3.2, oversupply and undersupply equilibria coexist. The undersupply equilibria are inefficient ones (sometimes referred to as WGC equilibria): comparing (21) and (22), the pickup time is longer in the undersupply equilibrium because a greater proportion of the mass is engaged in picking up customers (represented by the second term on the left-hand side of (22)). In such inefficient equilibria, drivers are assigned to pick up customers located far away and the system experiences a driver shortage because: 1) high average pick-up times force drivers to travel long distances to reach customers; 2) wasted time en route further reduces driver availability, exacerbating the undersupply. This creates a vicious cycle: longer travel times lead to fewer drivers being available to respond to new requests, worsening the imbalance between supply and demand. In the following, we show that theoretically the efficiency loss can be infinitely large.
@@
Infinity Price of Anarchy. The undersupply equilibrium can be arbitrarily worse compared to the oversupply equilibrium, i.e., $x^{u}<<b$. Specifically, when the mass is $m=\tfrac{1}{3}\left(\tfrac{b^{2}}{16v^{2}}\right)^{\frac{1}{3}}+bt$, there are exactly two equilibria. The two equilibria are $x^{o}=b$ and $x^{u}$, where $\frac{x^{u}}{b}$ satisfies
@@
which approaches 00 if $\theta\rightarrow\infty$. We summarize the results in the following corollary.
Corollary 3.3.
When $m=\tfrac{1}{3}\left(\tfrac{b^{2}}{16v^{2}}\right)^{\frac{1}{3}}+bt$ (case 2) of Proposition 3.2), the ratio between the two equilibrium driver supply rates, $x^{u}$ and $x^{o}$, tends to 0 as the waiting customers’ abandonment rate, $\theta$, goes to infinity, i.e.,
$\lim_{\theta\rightarrow\infty}\frac{x^{u}}{x^{o}}=0.$
Therefore, the inefficient equilibrium ($x^{u}$) can be arbitrarily worse than the efficient equilibrium ($x^{o}$). Note that an infinity price of anarchy occurs under the following conditions: 1) the customers exhibit extreme impatience ($\theta\rightarrow\infty$); 2) the region size is sufficiently large ($a\rightarrow\infty$ in order to apply the square root law (20) in this extreme case). This extreme scenario is primarily theoretical. However, it highlights a critical insight: if the dispatch algorithm is not properly designed, the system’s efficiency loss can be catastrophic.
@@
Figure 2. The LHS of (23) as a function of the driver supply rate $x$. We set the matching radius $R_{c}$ and $R_{d}$ to be the same. In this figure, the matching radius values, from left to right, are 1 km, 2 km, 3 km, and 4 km. We can observe that for larger matching radii ($R=2,3,4$ km) and a driver density near 25, there can be up to three equilibria. Additionally, the efficiency loss in the inefficient equilibrium is approximately $5\%-10\%$.
Impact of Matching Radius. In the following we investigate how different matching radii affect the outcome of equilibria and reveal the importance of the choice of matching radii. In the one region case, we can compute the equilibria by solving the equation of conservation law of mass in (3), i.e.,
@@
where the driver supply rate $x$ is the unknown and formulas of $w^{d}(x)$ and $\tau(x)$ are given by (16) and (17). Since $\tau(x)$ is non-monotonic, multiple equilibria can arise for specific parameter choices. Hence, the use of fixed matching radius with $R_{c},R_{d}<\sqrt{a/\pi}$, does not completely eliminate multiple equilibria or prevent the inefficient equilibrium.
@@
Extension to Multiple Region. When multiple regions are considered, the average pick-up time in each region becomes a non-monotonic function of the local driver supply rate, potentially resulting in a greater number of equilibria. For the single-region case, there may be up to three equilibria, but as the number of regions increases, the number of equilibria grows combinatorially. Due to the complexity of the detailed analysis, we omit it here, though this phenomenon can be verified using a simplified model, such as a spoke-and-hub network.
An important observation for systems with multiple regions is as follows: if the sum of the driver’s average waiting time, $w_{i}^{d}$, and the pick-up time, $\tau_{i}$ is an increasing function of the driver supply rate $x_{ii}$ for any region $i\in S$, then by Corollary 2.5, there is a unique equilibrium. However, under a general choice of matching radii $(R_{\cdot,c},R_{\cdot,d})$, $w_{i}^{d}+\tau_{i}$ is not necessarily an increasing function of $x_{ii}$. In the next section, we address this challenge and obtain monotonicity by dynamically optimizing the matching radii.